OpenStudy (anonymous):

A straight line L passing through P(2,3) cuts the +ve x-axis and the +ve y-axis at A and B respectively. If PA=2PB, find (a) the coordinates of A and B. (b) the equation of L.

OpenStudy (anonymous):

PA = 2PB... what does this mean?

OpenStudy (anonymous):

We know the slow is negative, since it passes through the positive axis. We also know a point it passes. What we need is a bit more information about the slope...

OpenStudy (anonymous):



OpenStudy (anonymous):

If A = 2B, then B/A = 1/2, so our slope is -1/2 ....? Am I interpreting it correctly?

OpenStudy (anonymous):

I don't know actually...

OpenStudy (anonymous):

Is PA the distance between P and A?

OpenStudy (anonymous):

The distance is unknown. PA is just about the straight line.

OpenStudy (anonymous):

Does PA = 2PB mean the length of PA is twice that of PB?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah, okay. Well let's just throw out some facts... \(A = (a, 0)\)\[ |PA| = \sqrt{(a-2)^2 + 3^2} \\ |PB| = \sqrt{3^2 + (b-3)^2} \]So\[ \sqrt{(a-2)^2 + 3^2} = 2 \cdot \sqrt{3^2 + (b-3)^2} \]In additon: \[ \sqrt{(a-2)^2 + 3^2} + \sqrt{3^2 + (b-3)^2} = \sqrt{a^2+b^2} \]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Oops, fluttered up a bit: \[ |PB| = \sqrt{2^2+(b-3)^2} \]

OpenStudy (anonymous):

ok......

OpenStudy (anonymous):

I'm just using pythagorean theorem... does it make sense?

OpenStudy (anonymous):

however, to find out PB, we need to use distance formula..........

OpenStudy (anonymous):

\[ \sqrt{(a-2)^2 + 3^2} = 2 \cdot \sqrt{2^2 + (b-3)^2} \\ (a-2)^2 + 3^2 = 4 [ 2^2 + (b-3)^2] \\ (a-2)^2+9 = 16+ 4(b-3)^4 \]

OpenStudy (anonymous):

The distance formula comes from the Pythagorean theorem.